# Disentangling the influence of excitation energy and compound nucleus angular momentum on fission fragment angular momentum

### v2

@@ Abstract
We report on IYR of 17 isotopes produced in the 28 MeV $\alpha$-induced fission of 232Th at the IGISOL facility of the University of Jyväskylä. The fissioning nuclei in this reaction are 233,234,235U*. We compare our data to IYR from thermal neutron-induced fission of 233U and 235U, and we observe statistically significant larger IYR in the 232Th($\alpha$,f) reaction, where the average compound nucleus (CN) spin is 7.5 $\hbar$, than in 233,235U(nth,f), with average spins 2.5 and 3.5 $\hbar$, respectively.
To assess the influence of the excitation energy, we study literature data of IYR from photon-induced fission reactions, and find that the IYR are independent of the CN excitation energy. We conclude that the different IYR must be explained by the different CN spin alone. This implies that the FF angular momentum only partly comes from the fission process itself, and is in addition influenced by the angular momentum present in the CN.
@@
Any model describing angular momenta of FFs must account for the influence of both excitation energy and angular momentum of the compound nucleus (CN). The influence of CN conditions on FF spins has been the focus of extensive research, both through experimental activities [5, 19, 20] and theoretical modeling [21, 8, 22].
@@
From an experimental perspective, Bulgac’s result is supported by the work of Giha et al. [19] who observed an increase of the $\gamma$-ray multiplicity as a function of the kinetic energy of the incident neutrons in the fission reaction of 239Pu.
@@
In this work, we investigate how the CN state affects the angular momentum of FFs by making use of the IYR and the assumption of a positive correlation between the FFs spin and the population of high-spin isomeric states in the observed fission products. We define the IYR as
@@
According to optical model calculations with TALYS using default parameters, the CN produced by 28 MeV $\alpha$-particle impinging on 232Th has an angular momentum of approximately 7.5 $\hbar$. We assume that pre-scission neutrons do not, on average, change the spin of the system, an assumption commonly made in studies of de-exciting fission fragments. On the other hand, thermal (s-wave) neutrons produce 236U* and 234U* with average angular momenta of about 3.5 $\hbar$ and 2.5 $\hbar$, respectively. Therefore, the CN produced in these three fission reactions are similar but carry significantly different angular momenta.
@@
We measured 17 IYRs from the 232Th($\alpha$,f) reaction using the Phase-Imaging Ion-Cyclotron-Resonance (PI-ICR) technique of the JYFLTRAP double Penning trap mass spectrometer, at the IGISOL-4 facility [39, 40, 30]. The technique has previously been used to measure IYR [41, 42, 43] and is applicable for isomeric states with half-lives longer than approximately 500 ms, and can resolve states separated by 30-40 keV, or more. The obtained data is presented in Table 1. Out of these, 16 have been measured for this system for the first time.
@@
The measured IYR from the present study are listed in Table 1 and shown in Figure 3. Except for 132Sb, all nuclei show a dominant feeding of the high-spin state, i.e. an IYR larger than 0.5. Figure 3 also shows the comparison between the IYRs measured in this study and those measured in the neutron-induced fission reactions retrieved from literature data, where the EXFOR data are averaged into one data point for each nucleus and fissioning system. The 232Th IYR data consistently show larger values than the 233U and 235U data (Figure 3). The weighted means of the difference between experimental and literature data are $\Delta$IYR = 0.145(8) and 0.157(6) for 235U and 233U, respectively. Normalizing with the change in CN spin as described above (from 7.5 $\hbar$ to 3.5 and 2.5 $\hbar$), we arrive at $\Delta$IYR/$\Delta\overline{\text{J}}$ = 0.033(2) $\hbar^{-1}$ for the studied isomeric pairs. The Wilcoxon signed-rank test suggests a significant difference between our measurements and the literature values (p = 9$\cdot$10-4) for 235U, and a strong trend towards a difference (p = 6.25$\cdot$10-2) for 233U. These data show that the higher spin CNs produced by 232Th($\alpha$,f) tend to promote more frequent de-excitation to higher spin isomeric states.
@@
Figure 4 shows IYR for the nuclei in this study for photon-induced fission reactions, as a function of the incident photon energy for different fissioning systems. These bibliographic data are retrieved from EXFOR [34]. We obtain an average slope of $\Delta IYR/\Delta E=0.0001(5)MeV^{-1}$, i.e., we see no effect of the incident photon energy on the IYR. This implies that the angular momentum of the FFs remains unaffected by an increase in the CN excitation energy. Part of the additional excitation energy can be carried away by pre-scission neutrons. Therefore, this observation also implies that neutron emission does not, at least not on average, remove a significant amount of angular momentum. This agrees with the result by Gjestvang et al. [20], and the common assumption that statistical neutrons carry away a negligible amount of angular momentum. This has, however, recently been questioned by Stetcu et al. [51].
@@
This contradicts the results from the modeling of the momentum generation in FREYA [8], but agrees with the theoretical and experimental observations of Bulgac et al. [22] and Giha et al. [19].
In their research, Wilson et al. [5] studied neutron-induced fission of 232Th and 238U at about 2 MeV incident energy, as well as spontaneous fission of 252Cf. From their results they concluded that a FFs’ angular momentum does not depend on the fissioning system; however, the CN angular momenta in these cases are all close to zero. Combining their findings with our results implies that while a part of the FF angular momentum comes from the fission process itself, it is also influenced by the angular momentum present in the CN.
Based on our results, one can estimate a lower limit of the additional angular momentum of a FF that is due to an increase in CN angular momentum. Using the TALYS code, we have estimated the additional angular momentum that a FF needs to have to increase the observed IYR by 0.15. We found that, on average, this increase needs to be at least one unit of $\hbar$. This would imply that at least 2 $\hbar$ out of the additional 4-5 $\hbar$ (i.e., $>$40 %) goes to the FFs’ angular momenta, while a maximum of 3 $\hbar$ (i.e., $<$60%) is converted into orbital angular momentum or dissipated via emission of pre-fission neutrons. A more careful calculation could be done using methods outlined in Refs.[28, 31, 30]. The values given here should be considered as a conservative estimate of a lower limit.

### v4

@@ Abstract
We report on IYR of 17 isotopes produced in the 28 MeV $\alpha$-induced fission of 232Th at the IGISOL facility of the University of Jyväskylä. The fissioning nuclei in this reaction are 233,234,235U*. We compare our data to IYR from thermal neutron-induced fission of 233U and 235U, and we observe statistically significant larger IYR in the 232Th($\alpha$,f) reaction, where the average compound nucleus (CN) spin is 7.7 $\hbar$, than in 233,235U(nth,f), with average spins 2.6 and 3.6 $\hbar$, respectively.
To assess the influence of the excitation energy, we study literature data of IYR from photon-induced fission reactions, and find that, within current uncertainties, the IYR indicate no dependency of the CN excitation energy. We conclude that the different IYR seem to be due to the different CN spins alone. This would imply that the FF angular momentum only partly comes from the fission process itself, and is in addition influenced by the angular momentum present in the CN.
@@
Any model describing angular momenta of FFs must account for the influence of both excitation energy and angular momentum of the compound nucleus (CN). The influence of CN conditions on FF spins has been the focus of extensive research, both through experimental activities [5, 19, 20, 21, 22, 23] and theoretical modeling [24, 8, 25].
@@
From an experimental perspective, Bulgac’s result is supported by References [19, 22, 21] which report an increase of the $\gamma$-ray multiplicity as a function of the kinetic energy of the incident neutrons in the fission reactions of 238U and 239Pu.
@@
In this work, we investigate how the CN state affects the angular momentum of FFs by making use of the IYR and the assumption of a positive correlation between the FFs spin and the population of high-spin isomeric states in the observed fission products.
We define the IYR as
@@
Figure 2: Relative formation probabilities of CN angular momenta resulting from the three considered reactions. The average values are 2.6 $\hbar$ and 3.6 $\hbar$ for 234U* and 236U*, respectively, when produced from thermal neutrons, and 7.7 $\hbar$ for 236U* produced from incident alpha particles at 28 MeV.
According to optical model calculations with TALYS using default parameters, the CN produced by 28 MeV $\alpha$-particle impinging on 232Th has an average angular momentum of 7.7 $\hbar$, as shown in Figure 2. We assume that pre-scission neutrons do not, on average, change the spin of the system, an assumption commonly made in studies of de-exciting fission fragments. The thermal (s-wave) neutrons produce 236U* and 234U* with average angular momenta of about 3.6 $\hbar$ and 2.6 $\hbar$ (Fig. 2), respectively. Therefore, the CN produced in these three fission reactions are similar but carry significantly different angular momenta. We assume that these distributions represent the angular momenta of those nuclei that actually fission.
@@
We measured 17 IYRs from the 232Th($\alpha$,f) reaction using the Phase-Imaging Ion-Cyclotron-Resonance (PI-ICR) technique of the JYFLTRAP double Penning trap mass spectrometer, at the IGISOL-4 facility [42, 43, 33]. The technique has previously been used to measure IYR [32, 44, 45] and is applicable for isomeric states with half-lives longer than approximately 500 ms, and can resolve states separated by 30-40 keV, or more. The obtained data are presented in Table 1. Out of these, 16 have been measured for this system for the first time.
@@
The measured IYR from the present study are listed in Table 1 and shown in Figure 4. Except for 132Sb, all nuclei show a dominant feeding of the high-spin state, i.e. an IYR larger than 0.5. Figure 4 also shows the comparison between the IYRs measured in this study and those measured in the neutron-induced fission reactions retrieved from literature data, where the EXFOR data are averaged into one data point for each nucleus and fissioning system. The 232Th IYR data consistently show larger values than the 233U and 235U data (Figure 4). The weighted means of the difference between experimental and literature data are $\Delta$IYR = 0.135(8) and 0.156(6) for 235U and 233U, respectively. Normalizing with the change in CN spin as described above (from 7.7 $\hbar$ to 3.6 and 2.6 $\hbar$), we arrive at $\Delta$IYR/$\Delta\overline{\text{J}}$ = 0.032(1) $\hbar^{-1}$ for the studied isomeric pairs. The Wilcoxon signed-rank test suggests a significant difference between our measurements and the literature values (p = 9$\cdot$10-4) for 235U, and a strong trend towards a difference (p = 6.25$\cdot$10-2) for 233U. These data show that the higher spin CNs produced by 232Th($\alpha$,f) tend to promote more frequent de-excitation to higher spin isomeric states.
The used literature data have been obtained with techniques different from the one employed in our work, typically using $\gamma$-ray spectroscopy. If provided, we use reported independent IYR but chose to include also cumulative IYR in Figure 4 and the derived numerical values. We verified that a restriction to independent IYR gives, within the uncertainties, the same results. Direct ion-counting is a more advanced technique for measuring IYR (see, e.g., [32]), and the difference in measurement methodology may explain some of the observed differences. However, the found difference in $\Delta$IYR for the two thermal-fission cases as given above is 0.02(1), while directly comparing the two cases gives 0.03(1). Since the difference in average CN spin is 1 $\hbar$, these values are in line with the proposed trend. Nevertheless, the reported values can only serve as preliminary estimates of a IYR dependency on the angular momentum of the compound nucleus.
@@
Figure 5 shows IYR for the nuclei in this study for photon-induced fission reactions, as a function of the incident photon energy for different fissioning systems. These bibliographic data are retrieved from EXFOR [37]. We obtain an average slope of $\Delta IYR/\Delta E=0.0001(5)MeV^{-1}$, i.e., we see no effect of the incident photon energy on the IYR. This implies that the angular momenta of the FFs seem to remain unaffected by an increase in the CN excitation energy.
Part of the additional excitation energy can be carried away by neutrons before and after scission. Therefore, this observation also implies that neutron emission does not, at least not on average, remove a significant amount of angular momentum. This agrees with the result by Gjestvang et al. [20], and the common assumption that statistical neutrons carry away a negligible amount of angular momentum. This has, however, recently been questioned by Stetcu et al. [53] for post-scission neutrons.
@@
If confirmed, this would contradict the results from the modeling of the momentum generation in FREYA [8], but agrees with the theoretical and experimental observations of References [25, 19, 22, 21, 23]. A PI-ICR measurement of IYR from thermal-neutron induced fission of 233,235U would be highly desirable and can clarify the situation.
In their research, Wilson et al. [5] studied neutron-induced fission of 232Th and 238U at about 2 MeV incident energy, as well as spontaneous fission of 252Cf. From their results they concluded that a FFs’ angular momentum does not depend on the fissioning system. The same lack of dependency is observed by Sears et al. [40] in their IYR compilation. They show, for a number of thermal neutron-induced fission reactions, that the measured IYR do not depend on the mass of the fissioning nuclei. However, the considered reactions dominantly have small CN angular momenta. Combining these findings with our results would imply that while a part of the FF angular momentum comes from the fission process itself, it is also influenced by the angular momentum present in the CN.
Based on our results, one can estimate a lower limit of the additional angular momentum of a FF that is due to an increase in CN angular momentum. Using the TALYS code, we have estimated the additional angular momentum that a FF needs to have to increase the observed IYR by 0.15. We found that, on average, this increase needs to be at least one unit of $\hbar$. This would imply that at least 2 $\hbar$ out of the additional 4-5 $\hbar$ (i.e., $>$40 %) goes to the FFs’ angular momenta, while a maximum of 3 $\hbar$ (i.e., $<$60%) is converted into orbital angular momentum or dissipated via emission of pre-fission neutrons. A detailed calculation could be done using methods outlined in Refs.[31, 34, 33]. The values given here should be considered as a conservative estimate of a lower limit.